I have a question that states: *
express in simplest form: 6/3x-2+3/2x
How do I do these two problems?

Answer by stanbon(48535)

*given f(x)=8/x+5 and g(x)=6/x+7, find (f-g)(x).
(f-g)(x) means f(x)-g(x)
Simply subtract g(x) from f(x)
--------------------
*given f(x)=9-x^2, and g(x)=3-x, find f/g(x)
f/g(x) means [f(x)]/[g(x)]
Divide f(x) by g(x)
======================
Cheers,
Stan H..